There are four roads from city X to Y and five roads from city Y to Z, find


(i) how many ways is it possible to travel from city X to city Z via city Y.


(ii) how different round trip routes are there from city X to Y to Z to Y and back


to X.

1
Expert's answer
2021-12-22T16:07:55-0500

(i)

Number of ways from "X" to "Y" is "4."

Number of ways from "Y" to "Z" is "5."

Then number of ways from "X" to "Z" passing "Y" is "4\cdot 5=20."


(ii)

Number of ways from "X" to "Z" passing "Y" is "4\cdot 5=20."

Number of ways from "Z" to "Y" without using the same road more than once is "5-1=4."

Number of ways from "Y" to "X" without using the same road more than once is "4-1=3."

Number of ways from "Z" to "X" passing "Y" without using the same road more than once is "4\cdot 3=12."

Number of different round trip routes are there from city "X" to "Z" passing "Y" and back from city "Z" to "X" passing "Y" is

"20\cdot12=240."